OpenStudy (anonymous):

an underwater (but near the surface) explosions is detected by sonar on a ship 50s before it is heard on the deck. if sound travels at 8000ft/s in water and 1300ft/s in air, how far is the ship from the explosion?

OpenStudy (anonymous):

i haved=8000t,d=1300(t+50)=1300t+65000

OpenStudy (anonymous):

i'm not sure what to do next

OpenStudy (phi):

d= 1300t+65000 d =8000t you could subtract the 2 equations, and then solve for t once you find t, you can find d

OpenStudy (phi):

or you can set the d's equal (it is the same distance) d= 1300t+65000 d =8000t so you can say 8000t = 1300t+65000 now solve for t, and then use that t to find d

OpenStudy (anonymous):

t=650/67

OpenStudy (phi):

now find d

OpenStudy (anonymous):

77612